Fit a least square linear approximation Y=a+bX for the data X: -1, 0, 1, 2, 3 and Y: 0, 1, 2 , 9, 29  


1
Expert's answer
2022-04-11T16:30:01-0400

### Theory ###

The least square linear approximation finds a and b coefficients by minimizing the sum of squares of differences between "a+bx_i" and "y_i" :

"S=\sum_{i=1}^n(a+bx_i-y_i)^2=min\Rarr\\
\Rarr\frac{\partial S}{\partial a}=0,\space \frac{\partial S}{\partial b}=0\\
\frac{\partial S}{\partial a}=\sum_{i=1}^n2(a+bx_i-y_i)=\\
=2(an+b\sum_{i=1}^{n}x_i-\sum_{i=1}^{n}y_i)=0\Rarr\\
\Rarr a=\frac{1}{n}(\sum_{i=1}^{n} y_i-b\sum_{i=1}^{n}x_i)\\
\frac{\partial S}{\partial b}=\sum_{i=1}^{n}2(a+bx_i-y_i)x_i=\\
2(a\sum_{i=1}^{n}x_i+b\sum_{i=1}^{n}x_i^2-\sum_{i=1}^{n}x_iy_i)\Rarr\\
\Rarr [insert\space a=\frac{1}{n}(\sum_{i=1}^{n} y_i-b\sum_{i=1}^{n}x_i)]\Rarr\\
\Rarr \sum_{i=1}^{n}x_i\sum_{i=1}^{n}y_i+\\
+b[n\sum_{i=1}^{n}x_i^2-(\sum_{i=1}^{n}x_i)^2]-n\sum_{i=1}^{n}x_iy_i=0\Rarr\\
b=(n\sum_{i=1}^{n}x_iy_i-\sum_{i=1}^{n}x_i\sum_{i=1}^{n}y_i)/\\
/(n\sum_{i=1}^{n}x_i^2-(\sum_{i=1}^{n}x_i)^2)"


### Solution ###

"\sum_{i=1}^{n}x_i=-1+0+1+2+3=5\\
\sum_{i=1}^{n}x_i^2=1+0+1+4+9=15\\
\sum_{i=1}^{n}y_i=0+1+2+9+29=41\\
\sum_{i=1}^{n}x_iy_i=0+0+2+18+87=107\\
b=(5\cdot107-5\cdot41)/(5\cdot15-25)=6.6\\
a=(41-6.6\cdot5)/5=1.6\\
Answer:y=1.6+6.6x"